OpenStudy (anonymous):

HELP WITH POLYNOMIALS PLEASE!! This is a pre-calc project about designing a roller coaster, and I just can't figure it out! Here are the requirements: The amusement park you are designing for gave you the following coaster requirements: - your coaster ride must have at least 3 relative maxima and/or minima - the ride length must be at least 4 minutes - the coaster ride starts at 250 feet - the ride dives below the ground into a tunnel at least once. There must be at least one double root, 2 real roots, and imaginary roots. Help please!!!AND IT ALSO HAS TO GO BACK TO 250 FEET

OpenStudy (dumbcow):

ok from the requirements we know that polynomial must be at least a 6th degree double root -->2 , 2 real, 2 imaginary equal 6 total roots f(0) = 250 f(T) = 250 where T is time it takes to complete roller coaster, T>4 also the double root and 2 real roots must be within the interval 0<t<T a general format would look like this: \[f(t) = a(t-d)^2 (t-h)(t-k)\] --> a*d^2*h*k = 250

OpenStudy (dumbcow):

oops i forgot about the imaginary roots \[f(t) = a(t^2 +i)(t-d)^2 (t-h)(t-k)\]

OpenStudy (dumbcow):

because there are a lot of unknowns, i would pre-determine length of ride and real roots example: T = 5 h = 1 k = 2 d = 4 \[f(0) = a(0+i)(0-4)^2(0-1)(0-2) = 250\] \[f(5) = a(25+i)(5-4)^2 (5-1)(5-2) = 250\] \[\rightarrow 32 a i = 250\] \[\rightarrow 12a (25+i) = 250\] solve for a and i

OpenStudy (dumbcow):

\[a = \frac{25}{48}\] \[i = 15\]

OpenStudy (dumbcow):

@HelpMewithpolynomials